OpenStudy (anonymous):

Simplify (2)(5^x)+(3)(5^x)

OpenStudy (anonymous):

So. . 5^x (2 + 3)?

OpenStudy (anonymous):

guess so better known as \(5\times 5^x\) or even \(5^{x+1}\)

OpenStudy (anonymous):

Why 5^x+1? Why would you add the 1?

OpenStudy (anonymous):

because \[5\times 5^x=5^1\times 5^x\] add the exponents since the bases are the same and get \(5^{x+1}\)

OpenStudy (anonymous):

like if \(x=3\) you would have \[5\times 5^3=5^4\]

OpenStudy (anonymous):

oooh that makes sense. Thank you!